# Reassessing quantum-thermodynamic enhancements in continuous thermal machines

### v2

@@ Appendix B Full Counting Statistics
The Full Counting Statistics (FCS) formalism is used to compute the variances of the different input and output currents of the quantum thermal machines presented and their classical equivalent models. In this formalism a set of counting fields $\vec{\chi}:=\{\chi_{r}\}$ are introduced that keep track of the exchanges of energy quanta between the machine and the thermal reservoirs $r$. These lead to the derivation of a generalized master equation for an extended density operator $\rho_{G}(t,\{\chi_{r}\}$ ) depending on the fields. It reads [78]:
@@
In any case, in the limit $\{\chi_{r}\}\rightarrow 0$, we recover $\rho_{\mathrm{G}}(t)=\rho(t)$ and (B) reduces to the standard master equation (2) for the machine evolution. Moreover, as in the case of the original master equation, Eq. (B) can be linearized and written in the form
@@
$\mathcal{C}_{2}=-\dfrac{1}{a_{1}}\left(a_{0}^{\prime\prime}+a_{1}^{\prime}\mathcal{C}_{1}+a_{2}\mathcal{C}_{1}^{2}\right).$ (63)
@@
Replacing the expression for the cumulant $\mathcal{C}_{1}$ in Eqs. 62 into the above Eq. (63), we obtain the final expression for the variance of the probability current in the long-time limit:
${\rm Var}[\dot{N}]=-\frac{1}{a_{1}}[a_{0}^{{}^{\prime\prime}}-a_{1}^{{}^{\prime}}\Big{(}\frac{a_{0}^{{}^{\prime}}}{a_{1}}\Big{)}+a_{2}\Big{(}\frac{a_{0}^{{}^{\prime}}}{a_{1}}\Big{)}^{2}],$ (64)
which is used to obtain all the variances for heat and work currents [Eqs. (54) and (55)]
Appendix C Classical thermodynamic-equivalent models details
We start by obtaining the equations of motion for the density operator elements from the master equation in Lindblad form (2), from which we can obtain the expression for a generic element of the density operator:
$\begin{split}\dot{\rho}_{ij}=-\mathrm{i}\bra{i}[H(t),\rho]\ket{j}+\sum_{r}\sum_{k}\bra{i}\mathcal{D}_{k}^{(r)}[\rho]\ket{j}.\end{split}$ (65)
After expanding the form of the dissipators $\mathcal{D}_{k}^{(r)}[\rho]=L_{k}^{(r)}\rho L_{k}^{(r)\dagger}-\frac{1}{2}\{L_{k}^{(r)\dagger}L_{k}^{(r)},\rho\}$, we obtain the following terms contributing to the above expression:
$\begin{split}\bra{i}[H(t),\rho]\ket{j}=&(\epsilon_{j}-\epsilon_{i})\rho_{ij}+g(\delta_{i\mathrm{u}}\rho_{\mathrm{v}j}+\delta_{i\mathrm{v}}\rho_{\mathrm{u}j}\\ &-\delta_{\mathrm{v}j}\rho_{i\mathrm{u}}-\delta_{\mathrm{u}j}\rho_{i\mathrm{v}}),\\ \bra{i}L_{k}^{(r)}\rho L_{k}^{(r)\dagger}\ket{j}=&\sum_{n,m}\alpha_{ni}^{k}\alpha_{mj}^{k}\sqrt{\gamma_{ni}\gamma_{mj}}\rho_{nm},\\ \bra{i}L_{k}^{(r)\dagger}L_{k}^{(r)}\rho\ket{j}=&\sum_{n,m}\alpha_{im}^{k}\alpha_{nm}^{k}\sqrt{\gamma_{im}\gamma_{nm}}\rho_{nj},\\ \bra{i}\rho L_{k}^{(r)\dagger}L_{k}^{(r)}\ket{j}=&\sum_{n,m}\alpha_{nm}^{k}\alpha_{jm}^{k}\sqrt{\gamma_{nm}\gamma_{jm}}\rho_{in},\\ \end{split}$ (66)
where we can again use the fact that the Lindblad operators produce jumps only between levels with a fixed energy gap $\Delta\epsilon_{k}=\pm\Delta\epsilon_{r}$, determined by the reservoirs and rewrite the $\alpha$-terms as delta functions:
To specialize this result for the case of Hamiltonian-induced coherence we can use the fact that in (87) first expression the only non-zero part is the second term, since the first will vanish, because we can always go to the interaction picture and consider only the interaction term of the total Hamiltonian (or in the non-energetic case both interacting levels have the same energy). On the other hand, the $\delta$ functions in (67) can be simplified by taking into account the fact that in this case different transitions cannot have the same $\Delta\epsilon_{k}$ associated with them. Then (67) becomes:
$\begin{split}&\alpha_{ni}^{k}\alpha_{mj}^{k}=\delta(\Delta\epsilon_{in}-\Delta\epsilon_{k})\delta_{ij}\delta_{nm},\\ &\alpha_{im}^{k}\alpha_{nm}^{k}=\delta(\Delta\epsilon_{mi}-\Delta\epsilon_{k})\delta_{ni},\\ &\alpha_{nm}^{k}\alpha_{jm}^{k}=\delta(\Delta\epsilon_{mn}-\Delta\epsilon_{k})\delta_{nj}.\end{split}$ (68)
which results in the set of equations (23).
In the case of noise-induced coherence, we don’t have an interaction Hamiltonian, so the first expression on (87) doesn’t need to be considered. The $\delta$ functions in (67), on the other hand, have more terms, since there are now different transitions associated with the same $\Delta\epsilon_{k}$. Now (67) reads:
$\begin{split}\alpha_{ni}^{k}\alpha_{mj}^{k}=&\delta(\Delta\epsilon_{in}-\Delta\epsilon_{k})(\delta_{ij}\delta_{nm}+\delta_{i\mathrm{u}}\delta_{j\mathrm{v}}\delta_{nm}+\delta_{i\mathrm{v}}\delta_{j\mathrm{u}}\delta_{nm}\\ &+\delta_{n\mathrm{u}}\delta_{m\mathrm{v}}\delta_{ij}+\delta_{n\mathrm{v}}\delta_{m\mathrm{u}}\delta_{ij}),\\ \alpha_{im}^{k}\alpha_{nm}^{k}=&\delta(\Delta\epsilon_{mi}-\Delta\epsilon_{k})(\delta_{ni}+\delta_{i\mathrm{v}}\delta_{n\mathrm{u}}+\delta_{i\mathrm{u}}\delta_{n\mathrm{v}}),\\ \alpha_{nm}^{k}\alpha_{jm}^{k}=&\delta(\Delta\epsilon_{mn}-\Delta\epsilon_{k})(\delta_{nj}+\delta_{j\mathrm{v}}\delta_{n\mathrm{u}}+\delta_{j\mathrm{u}}\delta_{n\mathrm{v}}).\end{split}$ (69)
In view of (69) the equations of motion are:
for the levels $n\neq\{\text{u},\text{v}\}$, and:
$\begin{split}\dot{\rho}_{nn}=&\sum_{i}\left[\gamma_{in}\rho_{ii}-\left(\gamma_{ni}\rho_{nn}+\sqrt{\gamma_{\text{u}i}\gamma_{\text{v}i}}\text{Re}(\rho_{\text{uv}})\right)\right],\\ \dot{\rho}_{\text{uv}}=&\sum_{i}[2\sqrt{\gamma_{i\text{u}}\gamma_{i\text{v}}}\rho_{ii}-\sqrt{\gamma_{\text{u}i}\gamma_{\text{v}i}}(\rho_{\text{uu}}+\rho_{\text{vv}})\\ &-(\gamma_{\text{u}i}+\gamma_{\text{v}i})\rho_{\text{uv}}].\\ \end{split}$ (71)
Finally, we show that all energy currents in the quantum machine and the classical equivalent are equal, also in the case of noise-induced coherence. Taking into account the fact that the transition in contact with the “work” reservoir is not involved in the noise inducing mechanism (involving only transitions from or to the degenerate levels), we observe that the classical equivalent can mimic the current $\langle\dot{Q}_{\mathrm{w}}\rangle$ in the original machine if it has the same populations in the steady state (since Eq. (22) depends only on the populations of the density matrix). Then, using Eq. (16), we also have that
$\langle\dot{Q}_{\mathrm{w}}\rangle=\langle\dot{Q}_{\mathrm{w}}^{\mathrm{cl}}\rangle\implies\langle\dot{N}\rangle=\langle\dot{N}^{\mathrm{cl}}\rangle,$ (72)
$\langle\dot{N}^{\mathrm{cl}}\rangle$ being the corresponding probability current in the classical equivalent machine. The above relation stating the equivalence of the steady-state probability currents in quantum and classical engines implies, at the same time, that all the other energy currents in the original and classical equivalent machines are equal to their classical counterparts.
Appendix D Expressions from Inverse Counting Statistics
@@
In this appendix we provide the explicit form of the FCS matrix $W_{G}$ and the relevant coefficients of the characteristic polynomial used to obtain analytical expressions for the currents variances using the Inverse Counting Statistics method, as introduced in App. B. For convenience we rewrite here the expressions of the probability flux in the long-time limit in terms of the relevant coefficients for both the quantum machines and their classical-equivalent counterparts:
@@
In the following subsections we first provide the expressions for the general cases of $N$-level multi-cycle machines discussed in Sec. V for Hamiltonian-induced coherence and noise-induced coherence, together with their corresponding classical equivalents. Then we also include the particular expressions for the three quantum thermal machines used as main illustrative examples in Sec. VI. We stress that while the explicit final expressions of ${\rm Var}[\dot{N}]$ and ${\rm Var}[\dot{N}^{\rm cl}]$ as well as those of some of the coefficients are not included here for size reasons, all analytical expressions can be found in the supplemental GitHub repository [97].
D.1 Generic multi-cycle Hamiltonian-induced coherence machines
@@
Following the coarse-graining procedure described in Sec.V we obtain the FCS matrix $W_{G}(\chi)$ in Eq. (50) with a single counting field $\chi$ for quantum machines with Hamiltonian-induced coherence in one of their transitions. Recall that this matrix includes only the relation between the relevant coherences and the diagonal elements of the density matrix in the steady state, while those associated to coherences that become zero in the steady-state are not included. In the case of Hamiltonian-induced coherence this is the case of the coherent subspace (spanned by states |$\ket{u}$ and |$\ket{v}$), while coherences related to the monitoring state |$\ket{m}$ or the coarse-grained state |$\ket{S}$ (involving the rest of the system populations) are not included. Taking that into account, the matrix $W_{G}$ associated to the vector $\vec{p}_{G}(t)=(\rho_{\rm uu},\rho_{\rm vv},\rho_{\rm mm},\rho_{\rm ss},\mathrm{Im}[\rho_{\rm uv}])$ has the form
@@
$W_{G}(\chi)=\left(\begin{array}[]{ccccc}-\Gamma_{\text{vs}}&0&0&\Gamma_{\text{sv}}&-2g\\ 0&-\gamma_{\text{um}}-\Gamma_{\text{us}}&e^{-i\chi}\gamma_{\text{mu}}&\Gamma_{\text{su}}&2g\\ 0&e^{i\chi}\gamma_{\text{um}}&-\Gamma_{\text{ms}}-\gamma_{\text{mu}}&\Gamma_{\text{sm}}&0\\ \Gamma_{\text{vs}}&\Gamma_{\text{us}}&\Gamma_{\text{ms}}&-\Gamma_{\text{sm}}-\Gamma_{\text{su}}-\Gamma_{\text{sv}}&0\\ g&-g&0&0&-\Gamma\\ \end{array}\right),$ (74)
@@
where in the last row standing for the imaginary part of the coherence between states |$\ket{\rm u}$ and |$\ket{\rm v}$, we defined $\Gamma=(\Gamma_{\mathrm{vs}}+\Gamma_{\mathrm{us}}+\gamma_{\mathrm{um}})/2$. The most relevant inverse counting coefficients associated with this matrix are (the rest can be found in the supplemental repository[97]):
$\begin{split}&a_{0}^{\prime}=\gamma_{\text{mu}}\Gamma_{\text{sm}}\left(\Gamma_{\text{us}}\left(2g^{2}+\Gamma\Gamma_{\text{vs}}\right)+2g^{2}\Gamma_{\text{vs}}\right)-\Gamma_{\text{ms}}\gamma_{\text{um}}\left(2g^{2}\left(\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Gamma\Gamma_{\text{su}}\Gamma_{\text{vs}}\right),\\ &a_{0}^{\prime\prime}=\Gamma_{\text{ms}}\gamma_{\text{um}}\left(2g^{2}\left(\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Gamma\Gamma_{\text{su}}\Gamma_{\text{vs}}\right)+\gamma_{\text{mu}}\Gamma_{\text{sm}}\left(\Gamma_{\text{us}}\left(2g^{2}+\Gamma\Gamma_{\text{vs}}\right)+2g^{2}\Gamma_{\text{vs}}\right),\\ &a_{1}^{\prime}=-\left[\left(\Gamma_{\text{vs}}+\Gamma\right)\left(\Gamma_{\text{ms}}\Gamma_{\text{su}}\gamma_{\text{um}}-\gamma_{\text{mu}}\Gamma_{\text{sm}}\Gamma_{\text{us}}\right)\right].\\ \end{split}$ (75)
@@
$W_{G}^{\mathrm{cl}}(\chi)=\left(\begin{array}[]{cccc}-\gamma_{\text{vu}}^{\text{cl}}-\Gamma_{\text{vs}}&\gamma_{\text{vu}}^{\text{cl}}&0&\Gamma_{\text{sv}}\\ \gamma_{\text{vu}}^{\text{cl}}&-\gamma_{\text{vu}}{}^{\text{cl}}-\gamma_{\text{um}}-\Gamma_{\text{us}}&e^{-i\chi}\gamma_{\text{mu}}&\Gamma_{\text{su}}\\ 0&e^{i\chi}\gamma_{\text{um}}&-\Gamma_{\text{ms}}-\gamma_{\text{mu}}&\Gamma_{\text{sm}}\\ \Gamma_{\text{vs}}&\Gamma_{\text{us}}&\Gamma_{\text{ms}}&-\Gamma_{\text{sm}}-\Gamma_{\text{su}}-\Gamma_{\text{sv}}\\ \end{array}\right),$ (76)
@@
$\begin{split}&a_{0}^{\mathrm{cl}\,\prime}=\frac{\Gamma_{\text{ms}}\gamma_{\text{um}}\left(2g^{2}\left(\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Omega\Gamma_{\text{su}}\Gamma_{\text{vs}}\right)-\gamma_{\text{mu}}\Gamma_{\text{sm}}\left(\Gamma_{\text{us}}\left(2g^{2}+\Omega\Gamma_{\text{vs}}\right)+2g^{2}\Gamma_{\text{vs}}\right)}{\Gamma},\\ &a_{0}^{\mathrm{cl}\,\prime\prime}=-\frac{\Gamma_{\text{ms}}\gamma_{\text{um}}\left(2g^{2}\left(\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Omega\Gamma_{\text{su}}\Gamma_{\text{vs}}\right)+\gamma_{\text{mu}}\Gamma_{\text{sm}}\left(\Gamma_{\text{us}}\left(2g^{2}+\Omega\Gamma_{\text{vs}}\right)+2g^{2}\Gamma_{\text{vs}}\right)}{\Gamma},\\ &a_{1}^{\mathrm{cl}\,\prime}=\Gamma_{\text{ms}}\Gamma_{\text{su}}\gamma_{\text{um}}-\gamma_{\text{mu}}\Gamma_{\text{sm}}\Gamma_{\text{us}}.\\ \end{split}$ (77)
D.2 Three-level coherent amplifier
To obtain the form of $W_{G}$ for this case, we write the system of equations for all elements of $\rho_{G}$ given by (B), using the Hamiltonian and Lindblad operators given in Sec. II.1. We consider only the relevant elements of the matrix leading to non-zero values of the density operator $\pi$ in the steady state. That is, we include terms connecting the level populations and the imaginary part of the coherence between states |$\ket{0}$ and |$\ket{1}$ (see Appendix C). On the other hand, both the real part of $\pi_{12}$ and the real and imaginary part of the other coherences $\pi_{13}$ and $\pi_{23}$ become zero at the steady state and we don’t need to describe their evolution. The matrix reads:
$W_{G}=\left(\begin{array}[]{cccc}-\gamma_{02}&0&\gamma_{20}~{}\mathrm{e}^{-i\chi_{\mathrm{h}}}&2g\\ 0&-\gamma_{12}&\gamma_{21}~{}\mathrm{e}^{-i\chi_{\mathrm{c}}}&-2g\\ \gamma_{02}~{}\mathrm{e}^{i\chi_{\mathrm{h}}}&\gamma_{12}~{}\mathrm{e}^{i\chi_{\mathrm{c}}}&-(\gamma_{20}+\gamma_{21})&0\\ g&-g&0&-\dfrac{1}{2}(\gamma_{02}+\gamma_{01})\end{array}\right),$

### v4

@@ Appendix E Full Counting Statistics methods
The Full Counting Statistics (FCS) formalism is used to compute the variances of the different input and output currents of the quantum thermal machines presented and their classical equivalent models. In this formalism, a set of counting fields $\vec{\chi}:=\{\chi_{r}\}$ is introduced that keep track of the energy quanta exchanges between the machine and the thermal reservoirs $r$. These lead to the derivation of a generalized master equation for an extended density operator $\rho_{G}(t,\{\chi_{r}\}$ ) depending on the fields. It reads [89]:
@@
In any case, in the limit $\{\chi_{r}\}\rightarrow 0$, we recover $\rho_{\mathrm{G}}(t)=\rho(t)$ and (E) reduces to the standard master equation (2) for the machine evolution. Moreover, as in the case of the original master equation, Eq. (E) can be linearized and written in the form
@@
$\mathcal{C}_{2}=-\dfrac{1}{a_{1}}\left(a_{0}^{\prime\prime}+2a_{1}^{\prime}\mathcal{C}_{1}+2a_{2}\mathcal{C}_{1}^{2}\right).$ (72)
@@
Replacing the expression for the cumulant $\mathcal{C}_{1}$ in Eqs. (71) into the above Eq. (72), we obtain the final expression for the variance of the probability current in the long-time limit:
${\rm Var}[\dot{N}]=-\frac{1}{a_{1}}[a_{0}^{{}^{\prime\prime}}-2a_{1}^{{}^{\prime}}\Big{(}\frac{a_{0}^{{}^{\prime}}}{a_{1}}\Big{)}+2a_{2}\Big{(}\frac{a_{0}^{{}^{\prime}}}{a_{1}}\Big{)}^{2}],$ (73)
which is used to obtain all the variances for heat and work currents [Eqs. (63) and (64)].
Appendix F Expressions from Inverse Counting Statistics
@@
In this Appendix we provide the explicit form of the FCS matrix $W_{G}$ and the relevant coefficients of the characteristic polynomial used to obtain analytical expressions for the currents variances using the Inverse Counting Statistics method, as introduced in Appendix E. For convenience we rewrite here the expressions of the probability flux in the long-time limit in terms of the relevant coefficients for both the quantum machines and their classical-equivalent counterparts:
@@
In the following subsections we first provide the expressions for the general cases of $N$-level multi-cycle machines discussed in the Results section for Hamiltonian-induced coherence and noise-induced coherence, together with their corresponding classical equivalents. Then we also include the particular expressions for the three quantum thermal machines used as main illustrative examples. We stress that while the explicit final expressions of ${\rm Var}[\dot{N}]$ and ${\rm Var}[\dot{N}^{\rm cl}]$ as well as those of some of the coefficients are not included here for size reasons, all analytical expressions can be found in the supplemental GitHub repository [92].
F.1 Generic multi-cycle Hamiltonian-induced coherence machines
@@
Following the coarse-graining procedure described in the main text we obtain the FCS matrix $W_{G}(\chi)$ in Eq.  (59) with a single counting field $\chi$ for quantum machines with Hamiltonian-induced coherence in one of their transitions. Recall that this matrix includes only the relation between the relevant coherences and the diagonal elements of the density matrix in the steady state, while those associated to coherences that become zero in the steady-state are not included. In the case of Hamiltonian-induced coherence this is the case of the coherent subspace (spanned by states |$\ket{u}$ and |$\ket{v}$), while coherences related to the monitoring state |$\ket{m}$ or the coarse-grained state |$\ket{S}$ (involving the rest of the system populations) are not included. Taking that into account, the matrix $W_{G}$ associated to the vector $\vec{p}_{G}(t)=(\rho_{\rm uu},\rho_{\rm vv},\rho_{\rm mm},\rho_{\rm ss},\mathrm{Re}[\rho_{\rm{uv}}],\mathrm{Im}[\rho_{\rm uv}])$ has the form
@@
$W_{G}(\chi)=\left(\begin{array}[]{cccccc}-\gamma_{\text{vm}}-\Gamma_{\text{vs}}&0&\gamma_{\text{mv}}&\Gamma_{\text{sv}}&0&-2g\\ 0&-\gamma_{\text{um}}-\Gamma_{\text{us}}&e^{-i\chi}\gamma_{\text{mu}}&\Gamma_{\text{su}}&0&2g\\ \gamma_{\text{vm}}&e^{i\chi}\gamma_{\text{um}}&-\Gamma_{\text{ms}}-\gamma_{\text{mu}}-\gamma_{\text{mv}}&\Gamma_{\text{sm}}&0&0\\ \Gamma_{\text{vs}}&\Gamma_{\text{us}}&\Gamma_{\text{ms}}&-\Gamma_{\text{sm}}-\Gamma_{\text{su}}-\Gamma_{\text{sv}}&0&0\\ 0&0&0&0&-\Lambda&\Delta_{d}\\ g&-g&0&0&-\Delta_{d}&-\Lambda\\ \end{array}\right),$ (75)
@@
where in the last row standing for the imaginary part of the coherence between states |$\ket{\rm u}$ and |$\ket{\rm v}$, we defined $\Lambda=(\Gamma_{\mathrm{vs}}+\Gamma_{\mathrm{us}}+\gamma_{\rm{vm}}+\gamma_{\mathrm{um}})/2$. The most relevant inverse counting coefficients associated with this matrix are (the rest can be found in the supplemental repository[92]):
$a_{0}^{\prime}=$ $\Gamma_{\text{ms}}\gamma_{\text{um}}\left(\Gamma_{\text{su}}\left[\left(\Delta_{d}^{2}+\Lambda^{2}\right)\left(\gamma_{\text{vm}}+\Gamma_{\text{vs}}\right)+2g^{2}\Lambda\right]+2g^{2}\Lambda\Gamma_{\text{sv}}\right)+\gamma_{\text{mv}}\gamma_{\text{um}}(\Gamma_{\text{su}}\left[\Gamma_{\text{vs}}\left(\Delta_{d}^{2}+\Lambda^{2}\right)+2g^{2}\Lambda\right]$
$+2g^{2}\Lambda\Gamma_{\text{sm}}+2g^{2}\Lambda\Gamma_{\text{sv}})-\gamma_{\text{mu}}[\gamma_{\text{vm}}\left[\Delta_{d}^{2}\Gamma_{\text{us}}\left(\Gamma_{\text{sm}}+\Gamma_{\text{sv}}\right)+\Lambda\left(2g^{2}\left(\Gamma_{\text{sm}}+\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Lambda\Gamma_{\text{us}}\left(\Gamma_{\text{sm}}+\Gamma_{\text{sv}}\right)\right)\right]$
$+\Gamma_{\text{sm}}\left(\Gamma_{\text{us}}\left(\Gamma_{\text{vs}}\left(\Delta_{d}^{2}+\Lambda^{2}\right)+2g^{2}\Lambda\right)+2g^{2}\Lambda\Gamma_{\text{vs}}\right)],$ (76)
$a_{0}^{\prime\prime}=$ $-\Gamma_{\text{ms}}\gamma_{\text{um}}\left(\Gamma_{\text{su}}\left[\left(\Delta_{d}^{2}+\Lambda^{2}\right)\left(\gamma_{\text{vm}}+\Gamma_{\text{vs}}\right)+2g^{2}\Lambda\right]+2g^{2}\Lambda\Gamma_{\text{sv}}\right)-\gamma_{\text{mv}}\gamma_{\text{um}}(\Gamma_{\text{su}}\left[\Gamma_{\text{vs}}\left(\Delta_{d}^{2}+\Lambda^{2}\right)+2g^{2}\Lambda\right]$
$+2g^{2}\Lambda\Gamma_{\text{sm}}+2g^{2}\Lambda\Gamma_{\text{sv}})-\gamma_{\text{mu}}[\gamma_{\text{vm}}\left(\Delta_{d}^{2}\Gamma_{\text{us}}\left(\Gamma_{\text{sm}}+\Gamma_{\text{sv}}\right)+\Lambda\left[2g^{2}\left(\Gamma_{\text{sm}}+\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Lambda\Gamma_{\text{us}}\left(\Gamma_{\text{sm}}+\Gamma_{\text{sv}}\right)\right]\right)$
$+\Gamma_{\text{sm}}\left(\Gamma_{\text{us}}\left(\Gamma_{\text{vs}}\left(\Delta_{d}^{2}+\Lambda^{2}\right)+2g^{2}\Lambda\right)+2g^{2}\Lambda\Gamma_{\text{vs}}\right)],$
$a_{1}^{\prime}=$ $2\gamma_{\text{mv}}\gamma_{\text{um}}\left(g^{2}\left(\Lambda+\Gamma_{\text{sm}}+\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Lambda\Gamma_{\text{su}}\Gamma_{\text{vs}}\right)+\Gamma_{\text{ms}}\gamma_{\text{um}}\left(\Gamma_{\text{su}}\left[\Delta_{d}^{2}+2g^{2}+\Lambda^{2}+2\Lambda\left(\gamma_{\text{vm}}+\Gamma_{\text{vs}}\right)\right]+2g^{2}\Gamma_{\text{sv}}\right)$
$-\gamma_{\text{mu}}\left(\Gamma_{\text{sm}}\left(\Gamma_{\text{us}}\left(\Delta_{d}^{2}+2g^{2}+\Lambda^{2}+2\Lambda\Gamma_{\text{vs}}\right)+2g^{2}\Gamma_{\text{vs}}\right)+2\gamma_{\text{vm}}\left[g^{2}\left(\Lambda+\Gamma_{\text{sm}}+\Gamma_{\text{su}}+\Gamma_{\text{sv}}\right)+\Lambda\Gamma_{\text{us}}\left(\Gamma_{\text{sm}}+\Gamma_{\text{sv}}\right)\right]\right).$
@@
$W_{G}^{\mathrm{cl}}(\chi)=\left(\begin{array}[]{cccc}-\gamma_{\text{vu}}{}^{\text{cl}}-\gamma_{\text{vm}}-\Gamma_{\text{vs}}&\gamma_{\text{vu}}{}^{\text{cl}}&\gamma_{\text{mv}}&\Gamma_{\text{sv}}\\ \gamma_{\text{vu}}{}^{\text{cl}}&-\gamma_{\text{vu}}{}^{\text{cl}}-\gamma_{\text{um}}-\Gamma_{\text{us}}&e^{-i\chi}\gamma_{\text{mu}}&\Gamma_{\text{su}}\\ \gamma_{\text{vm}}&e^{i\chi}\gamma_{\text{um}}&-\Gamma_{\text{ms}}-\gamma_{\text{mu}}-\gamma_{\text{mv}}&\Gamma_{\text{sm}}\\ \Gamma_{\text{vs}}&\Gamma_{\text{us}}&\Gamma_{\text{ms}}&-\Gamma_{\text{sm}}-\Gamma_{\text{su}}-\Gamma_{\text{sv}}\\ \end{array}\right),$ (77)
@@
$\begin{split}&a_{0}^{\mathrm{cl}\,\prime}=\dfrac{a_{0}^{\prime}}{\Lambda^{2}+\Delta_{\rm d}^{2}},~{}~{}~{}~{}a_{0}^{\mathrm{cl}\,\prime\prime}=\dfrac{a_{0}^{\prime\prime}}{\Lambda^{2}+\Delta_{\rm d}^{2}},~{}~{}~{}~{}a_{1}^{\mathrm{cl}\,\prime}=\frac{2g^{2}\Lambda\left(\gamma_{\text{mv}}\gamma_{\text{um}}-\gamma_{\text{mu}}\gamma_{\text{vm}}\right)}{\Delta_{d}^{2}+\Lambda^{2}}+\Gamma_{\text{ms}}\Gamma_{\text{su}}\gamma_{\text{um}}-\gamma_{\text{mu}}\Gamma_{\text{sm}}\Gamma_{\text{us}}.\\ \end{split}$ (78)
F.2 Three-level coherent amplifier
To obtain the form of $W_{G}$ for this case, we write the system of equations for all elements of $\rho_{G}$ given by (E), using the Hamiltonian and Lindblad operators given in Eq. (26) and (27). We consider only the relevant elements of the matrix leading to non-zero values of the density operator $\pi$ in the steady state. That is, we include terms connecting the level populations and the imaginary part of the coherence between states |$\ket{0}$ and |$\ket{1}$ (see App. D). On the other hand, both the real part of $\pi_{12}$ and the real and imaginary part of the other coherences $\pi_{13}$ and $\pi_{23}$ become zero at the steady state and we don’t need to describe their evolution. The matrix reads:
$W_{G}=\left(\begin{array}[]{cccc}-\gamma_{02}&0&\gamma_{20}&2g\\ 0&-\gamma_{12}&\gamma_{21}~{}\mathrm{e}^{-i\chi}&-2g\\ \gamma_{02}&\gamma_{12}~{}\mathrm{e}^{i\chi}&-(\gamma_{20}+\gamma_{21})&0\\ g&-g&0&-\dfrac{1}{2}(\gamma_{02}+\gamma_{12})\end{array}\right),$